Find the slope-intercept equation of the line that passes through (1,3) and (3,-3).
Please show the work so I can see how the problem was worked out. Thanks

Answer by rothauserc(4718)

The standard formula for a line is y = mx +b, where m is the slope and b is the y intercept
m = (y2 - y1) / (x2 -x1) = (-3 -3) / (3 - 1) = -6/2 = -3
so far we have
y = -3x +b
to calculate b, substitute for x and y using one of the points, let's pick (1, 3)
3 = -3*(1) +b
3 = -3 + b
add 3 to both sides of =
b = 6
now we have the complete formula
y = -3x +6